20250221
Time
Attendance
Dimitry, Sarah, Saisri, Nick, Bhanu
Spaces Occupied
433, dome
Goals
astropol session
Equipment Used
Scope | Camera | Focus |
---|---|---|
Main | PICSI2 w/ 532MM | |
Finder | ASI 482 | 2465 |
Frame Table
Time Observed | Source Name | Filter | Int Time [s] | Gain | Airmass | Mag (filter) | RA [h] | DEC [deg] | Source Type | Scope | Purpose | Notes |
---|---|---|---|---|---|---|---|---|---|---|---|---|
YYYYmmddTHHMM | Object | Filter | time | gain | airmass | mag(V) | HH | DD | type | Scope | Purpose | tracking? Others? |
Log
- The south mirror door was the only one to open initially. Briefly jiggling the north door motor opened it and the other doors followed.
- Auto dome was not working correctly at first but worked after slewing to zenith and restarting DFMTCS
- Pointing took hours. Started around 1530, attempted to locate Venus and Jupiter but never found either. Managed to point using astrometry at around 1820 and successfully located Venus.
- Mirror door issue persisted after changing the instrument and while shutting down. Eileen's note: Issue was the East door being lapped on top of North. We eventually closed the dome to warm things back up, and about 15 minutes after that, I was trying to hold the limit switches on north and west while opening, to see if that would prevent east getting “bound” by those (stuck) doors before it could open. It did allow us to finally get them unstuck but it's not clear if it was the limit switch trick or just warming up the motors. Might consider installing small heating elements around the motors that can be turned on as needed if thermal state is your issue (this is actually not uncommon at other observatories I know of).
ToDo
Notes
- Note for Astropol: Focus 2015.0 SPRITE
Weather
1617 - 35 F, 41% humidity
1820 - 32 F, 45% humidity
2015 - 29 F, 53% humidity